Problem 5 (7 marks)

A survey of 36 retail stores revealed that the average price of a tablet was $375 with a standard deviation of $20. a) What is the 95% confidence interval to estimate the true cost of the tablet? (4 marks) b) What sample size would be needed to estimate the true average price of a tablet with an error of 5$ ± and a 99% confidence? (3 marks)
Answer #1

Solution :

Given that,

a) Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha/2 * ( \sigma /\sqrtn)

= 1.96 * ( 20 /  \sqrt36 )

= 6.53

At 95% confidence interval estimate of the population mean is,

\bar x  ± E

= 375  ± 6.53

= ( $368.47, $381.53 )

b) Z\alpha/2 = Z0.005 = 2.576

sample size = n = [Z\alpha/2* \sigma / E] 2

n = [ 2.576 * 20 / 5 ]2

n = 106.17

Sample size = n = 107